OpenStudy (scarlettfarra2000):

Could someone please see if I'm doing this right? I think I did the last part wrong Use the zero-product property to solve the following equation

OpenStudy (scarlettfarra2000):

The zero – product states that for any real numbers a and b, if ab=0 to use the zero – product property to solve equation, set each factor equal to zero and solve. First we set the factor -3n equal to zero and solve for n. -3n =0 Now we set the factors 2n-7 equal to zero and solve for n. Add 1 to get side and simplify Divide each side by 2 2n-7=0 -2n -2n

OpenStudy (atrineas):

Ok so 2n-7=0 first add 7 to both sides like so. . . 2n=7 then divide both sides by 2 like so. . . 2n/2=7/2 n=7/2

OpenStudy (anthonyym):

The first factor -3n. -3n = 0 n = 0, reasoning: divide both sides by -3, or 0 product property. Second factor 2n-7. 2n - 7 = 0 You are solving for n, that means to get n on by itself on one side of the equation. Add 7 to both sides to get -7 to the other side. 2n - 7 = 0 2n = 7 Then divide both sides by 2. n = 7/2

OpenStudy (scarlettfarra2000):

Oh I see thank you @anthonyym and @Atrineas

OpenStudy (anthonyym):

no problem

OpenStudy (atrineas):

^~^ no problem @ScarlettFarra2000 *High fives @anthonyym*

OpenStudy (anthonyym):

Lol

OpenStudy (atrineas):

xD